A mass m at the end of a spring vibrates with a frequency of 0.90 Hz. When an additional 520-g mass is added to m, the frequency is 0.50 Hz.  

   a) What is the value of m?                                                                                                

  b) Calculate the total mass needed for the system to vibrate with a freq. of 0.45 Hz. 


1
Expert's answer
2021-04-30T11:28:06-0400

"f = \dfrac{1}{2π}\sqrt{\dfrac{k}{m}}"


"k =4\u03c0\u00b2mf\u00b2"


"\therefore 4π²m_1f_1² = 4π²m_2f_2²"

"m_1f_1\u00b2 = m_2f_2\u00b2"


a.

"m_1\u00d7 0.90\u00b2 = (m_1+520) \u00d7 0.50\u00b2"

"0.90\u00b2 m_1= 0.50\u00b2m_1 +130"

"0.56m_1 = 130"

"m_1 =232.14g"


b.

"k =4\u03c0\u00b2mf\u00b2"

"k = 4\u03c0\u00b2\u00d7232.14 \u00d70.90 = 8248.07Nm^{-1}"


"m= \dfrac{k}{4π²f²} =\dfrac{8248.07}{4π²×0.45²} ="


"1031.73g \approx1.031kg"